You purchased a bond for $1,000. At the end of the year, the bond paid interest of $150. You sold the bond for $750 after receiving the interest payment. What was your total rate of return on this investment? What was your capital gains yield? What was your coupon yield? Please show work
Answer #1

Total rate of return=(End value-Beginning value+Interest)/Beginning value

=(750-1000+150)/1000

=(10%)(Negative).

Capital gains yield=(End value-Beginning value)/Beginning value

=(750-1000)/1000

=(25%)(Negative).

Coupon yield=Interest/Beginning value

=(150/1000)

=15%